OpenStudy (anonymous):

I have a business Math ques. I did the math but want to double check my ans: ---you deposit $1000 for 4 yrs at an int rate of 2.0%. if the int is compounded annually, how much money do you have after the 4 yrs? A) $1082.40; B) $1100; C) $1103.80; D) $1344.90 Also, same ques but int is compounded quarterly. A) $1081.70; B) $1083.10; C) $1110.80; D) $1351.30

OpenStudy (tkhunny):

How do you achieve those results? Let's see your formulation.

OpenStudy (anonymous):

ok. give me a min to go over my notes

OpenStudy (anonymous):

1st ques: i used an int table and got 1.0824 which i multiplied by $1000 and got $1082.40

OpenStudy (anonymous):

2nd ques: annual rate/# of qtrs 2/4=0.5 and then i used the same int table to find my decimal equivalent. 1.0831 x $1000 and got $1083.10

OpenStudy (tkhunny):

That explains why the results were rounded oddly. \(1000\cdot(1 + 0.02)^{4} = 1000\cdot(1.02)^{4} = 1000\cdot 1.08243216 = 1082.43216\), so a reasonable answer might be $1,082.43. \(1000\cdot(1 + 0.02/4)^{14} = 1000\cdot(1.005)^{16} = 1000\cdot 1.08307115 = 1083.07115\), so a reasonable answer might be $1,083.07.

OpenStudy (anonymous):

OMG, i got it right! thank you. u get a medal

OpenStudy (tkhunny):

Watch for that rounding stuff. It can be a bit annoying.

OpenStudy (anonymous):

thanks for the tip